An ACL patient is doing leg extensions to strength
Answer #1

Solution: 5 cm 7259 32 cm W = mg (a). Moment at hinge is. Em=0 F, sin 10°(5) -W (32)cos 25° = 0 F, sin 10°(5) =W (32)cos 25°R-F, cos 10°sin 65° + Fosin 10° sin 25º = 0 R = Focos 10°sin 65°- Fosin 10°sin 25° = (2045.93 N) cos 10° sin 65°-(2045.93 N)